Question1

 

Given   matrix A =        2      2     1
                          1      3     1
                          1      2     2

 
 

1.1 find the eigenvalues of   A

 
1.2.   Determine an eigenvector for the eigenvalue  =  5






1
Expert's answer
2021-08-05T08:02:12-0400

"A=\begin{bmatrix}
 2 & 2&1 \\
 1 & 3&1\\
1&2&2
\end{bmatrix}"

1.1. "\det (A-\lambda I)=\begin{vmatrix}
 2-\lambda & 2&1 \\
 1 & 3-\lambda &1\\ 1&2&2-\lambda 
\end{vmatrix}=-\lambda^3+7\lambda^2-11\lambda+5=-(\lambda-5)(\lambda-1)^2=0"


Eigenvalues: "\lambda _1=5" and "\lambda_2=1"


1.2.

"\lambda_1=5" ": \ \ (A-5I)\mathbf{x}= \begin{bmatrix}
 -3 & 2&1 \\
 1 & -2&1\\
1&2&-3
\end{bmatrix}
\begin{bmatrix}
 x\\y\\z
\end{bmatrix}=\begin{bmatrix}
 0\\0\\0
\end{bmatrix}" gives the eigenvector "\mathbf{x}_1=\begin{bmatrix}1\\1\\1\end{bmatrix}" .


Answer: 1.1. "\lambda_1=5" and "\lambda_2=1" ; 1.2. "\mathbf{x}_1=\begin{bmatrix}1\\1\\1\end{bmatrix}"